OpenStudy (anonymous):

Write an equation of the line that passes through the point (-4,1) and is (a) parallel to and (b) perpendicular to the line 2x + 7y=14

OpenStudy (anonymous):

(a)2x+7y=-1 (b)7x-2y=-30

OpenStudy (jamesj):

So two lines are parallel if they have the same slope; and perpendicular is their slopes are the negative reciprocal of each other. So the first step is the convert the equation 2x + 7y = 14 into a form where you can see the slope, call it m. Then for part (a) you now know the slope, m, and one point and you can write down the equation of the line. And for part (b) the same is also true with a slope -1/m.

OpenStudy (anonymous):

@jamesj, can u help me in my problem??